air leallon hat is looie nida upis positive. he coin attains Dr he maximuna bl find it aleve theo 0Ds attrbting released me sccood L.DDs after be ased Find tte time , in seconds, that it takes the ca to nit the
Answer #1

a) from kinematic equation y vo? = saya at maximum height, v= m/s - Vo = zay = 4.6om 2(-9-80m/5) maxincem height = 310m + 4-6